An equation of the line in slope intercept form that passes through the point (-9, -1) and is perpendicular to the line y = 3x + 17?
Answer #1
the slope you want is -1/3

Now you have a point and a slope:

(y- (-1))/(x - (-9)) = -1/3

or

(y+1)/(x+9) = -1/3

y+1 = -1/3 x - 3

y = -1/3 x - 4
or
x + 3y + 12 = 0
answered by: arbee